OpenStudy (anonymous):

Help Stephanie write one radical equation where the solution is extraneous and another equation where the solution is non-extraneous. Using complete sentences, explain each step when solving to justify your examples.

OpenStudy (anonymous):

@Yuba @HelpBlahBlahBlah @amoodarya @Awesome781 @denizen @micahpkay

OpenStudy (anonymous):

@Preetha @ParthKohli @waleedimtiaz @mustafa2014

OpenStudy (anonymous):

@gracerigg @sourwing @zobya @zimmah

OpenStudy (anonymous):

somebody please help me out

OpenStudy (anonymous):

@TuringTest @TheKylieeM

OpenStudy (turingtest):

a non-extraneous what? what math is this?

OpenStudy (anonymous):

algebra 2

OpenStudy (anonymous):

and a non-extraneous equation

OpenStudy (turingtest):

I had to look it up. You are talking about non-extraneous *solutions*, which is one that satisfies the original equation.

OpenStudy (anonymous):

https://www.youtube.com/watch?v=711pdW8TbbY

OpenStudy (anonymous):

yeaa

OpenStudy (turingtest):

your question is very poorly-worded; you want the "radical of a non-extraneous solution"? I don't understand what you want, and since you don't even seem to understand the term "extraneous solution", I'm guessing you don't know either. Do you have the exact wording of the question in English?

OpenStudy (anonymous):

Help Stephanie write one radical equation where the solution is extraneous and another equation where the solution is non-extraneous. Using complete sentences, explain each step when solving to justify your examples

OpenStudy (turingtest):

in the future, please put the question exactly in your initial post, you could probably have been helped much earlier

OpenStudy (turingtest):

So write an equation involving a \(\sqrt x\) somewhere, solve it, then check to see if it's a solution to the initial problem by plugging it back into the beginning. if it is, then it's non-extraneous.

OpenStudy (anonymous):

so is this right sqrt x = 2

OpenStudy (turingtest):

sure, now solve it

OpenStudy (anonymous):

ummm x=4 right

OpenStudy (turingtest):

right, now check to see that it solves the original equation

OpenStudy (anonymous):

yes it does

OpenStudy (turingtest):

so then it must be non-extraneous

OpenStudy (anonymous):

ok and for an extraneous solution would this work sqrt x=2x-12

OpenStudy (turingtest):

I don;t know, why don't you try it and show me?

OpenStudy (anonymous):

idk how to reaally do it what i got so far is x=4x^2-48x+24 then you have to subtract the x from both side which gives you 4x^2-49x+24

OpenStudy (anonymous):

@zimmah i looked at the youtube video but im stil confused

OpenStudy (anonymous):

do you know that quadratic equation

OpenStudy (anonymous):

i think thats the next step

OpenStudy (turingtest):

\[ax^2+bx+c=0\implies x={-b\pm\sqrt{b^2-4ac}\over 2a}\]

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

so then its -49 sqrt 2017 over 8

OpenStudy (anonymous):

@zimmah @TuringTest am i right so far

OpenStudy (turingtest):

remember to write the +/-

OpenStudy (turingtest):

and I don't feel like breaking out a calculator to solve the problem this way, so you;ll have to trust your own. just put the numbers in the right places.

OpenStudy (anonymous):

i'm not sure how you arrived at that answer, but it doesn't seem to be correct. Note how you can only use the quadratic formula if on the one side of the equals sign you have 0, and on the other side you have ax^2+bx+c